HQ7.38 Homework Unanswered Which of the following is not isoelectronic with the others? Cl A C Ar O TI+ E Rb
Answer #1

Isoelectronic species are those which contain the same number of electrons.

A. Atomic number of Chlorine (Cl) is 17. As number of electrons in a neutral atom is equal to the atomic number of the element so number of electrons in Cl are 17 . Negative charge of 1 unit on Cl means Cl has acquired an extra electron so Cl​​​​​- has 18 electrons in it.

Thus number of electrons in Cl​​​​​- = 18

B. Atomic number of sulfur (S) is 16. Thus number of electrons in S are 16 . Negative charge of 2 units on S means 2 electrons are gained by neutral S atom. Thus S​​​​​​2- has 16+2 = 18 electrons.

Thus number of electrons in S​​​​​​2- = 18

C. Atomic number of Argon (Ar) is 18. So number of electrons in neutral Ar atom are 18.

Thus number of electrons in Ar = 18

D. Atomic number of Titanium (Ti) is 22. Thus number of electrons in Ti are 22. Positive charge of 4 units on Ti means loss of 4 electrons by Titanium. So Ti​​​​​4+ has 22 - 4 = 18 electrons.

Thus number of electrons in Ti​​​​​4+ = 18

E. Atomic number of Rubidium (Rb) is 37. Thus number of electrons in Rb are 37. Positive charge of 1 unit on Rb means loss of one electron by Rubidium. So Rb​​​+ has 37-1 = 36 electrons.

Thus number of electrons in Rb​​​+ = 36

Since Cl​​​​​- , S​​​​​​2- , Ar and Ti​​​​​4+ all have 18 electrons so they are isoelectronic. Rb​​​+ having 36 electrons is not isoelectronic with the others.

Hence option E is the correct answer.